the perimeter of a rectangle is 54 m. The width is 3 m less than twice the length. Find the dimensions of the rectangle?
Answer by JBarnum(2146)

GIVEN:
2W+2L=P
P=54
W=2L-3
,
SOLVE:
substitution
2(2L-3)+2L=54
4L-6+2L=54
6L=60
L=10
.
W=2L-3
W=2(10)-3
W=20-3
W=17